# AgentClinic: a multimodal agent benchmark to evaluate AI in simulated clinical environments

### v1

@@ Abstract
Diagnosing and managing a patient is a complex, sequential decision making process that requires physicians to obtain information—such as which tests to perform—and to act upon it. Recent advances in artificial intelligence (AI) and large language models (LLMs) promise to profoundly impact clinical care. However, current evaluation schemes overrely on static medical question-answering benchmarks, falling short on interactive decision-making that is required in real-life clinical work. Here, we present AgentClinic: a multimodal benchmark to evaluate LLMs in their ability to operate as agents in simulated clinical environments. In our benchmark, the doctor agent must uncover the patient’s diagnosis through dialogue and active data collection. We present two open benchmarks: a multimodal image and dialogue environment, AgentClinic-NEJM, and a dialogue-only environment, AgentClinic-MedQA. We embed cognitive and implicit biases both in patient and doctor agents to emulate realistic interactions between biased agents. We find that introducing bias leads to large reductions in diagnostic accuracy of the doctor agents, as well as reduced compliance, confidence, and follow-up consultation willingness in patient agents. Evaluating a suite of state-of-the-art LLMs, we find that several models that excel in benchmarks like MedQA are performing poorly in AgentClinic-MedQA. We find that the LLM used in the patient agent is an important factor for performance in the AgentClinic benchmark. We show that both having limited interactions as well as too many interaction reduces diagnostic accuracy in doctor agents. The code and data for this work is publicly available at AgentClinic.github.io.
@@ Introduction
    Results of the diagnostic accuracy of four language models on AgentClinic-MedQA: GPT-4, Mixtral-8x7B, GPT-3.5, and Llama 2 70B-chat. We also evaluate two language models on the multimodal AgentClinic-NEJM benchmark: GPT-4-turbo and GPT-4-vision-preview.
@@ Results
Here we discuss the accuracy of various language models on AgentClinic-MedQA. We evaluate four models in total: GPT-4, Mixtral-8x7B, GPT-3.5, and Llama 2 70B-chat (discussed in detail in Appendix A). Each model acts as the doctor agent, attempting to diagnose the patient agent through dialogue. The doctor agent is allowed N=20 patient and measurement interactions before a diagnosis must be made. For this evaluation, we use GPT-4 as the patient agent for consistency. The accuracies of the models are presented in Figure 5: GPT-4 at 52%, GPT-3.5 at 38%, Mixtral-8x7B at 37%, and Llama 2 at 70B-chat 9%.
@@
Many types of diagnoses require the physician to visually inspect the patient, such as with infections and rashes. Additionally, imaging tools such as X-ray, CT, and MRI provide a detailed and rich view into the patient, with hospitalized patients receiving an average of 1.42 diagnostic images per patient stay[25]. However, the previous experiments in this work and prior work[26] provided measurement results through text, and did not explore the ability of the model to understand visual context. Here, we evaluate two multimodal LLMs, GPT-4-turbo and GPT-4-vision-preview, in a diagnostic settings that require interacting through both dialogue as well as understanding image readings. We collect our questions from New England Journal of Medicine (NEJM) case challenges. These published cases are presented as diagnostic challenges from real medical scenarios, and have an associated pathology-confirmed diagnosis. We curate 15 challenges from a sample of 932 total cases for AgentClinic-NEJM. While for human viewers, these cases are provided with a set of multiple choice answers, we chose to not provide these options to the doctor agent and instead keep the problems open-ended.
@@
When the image is provided initially to the doctor agent, across 15 multimodal patient settings we find that GPT-4-turbo and GPT-4-vision-preview obtain an accuracy of 27% (Fig. 8). We also find that for the provided incorrect responses from GPT-4-turbo, the answer that was provided was among those listed in the multiple choice options 60% of the time. Despite having the same accuracy, we find that GPT-4-vision-preview was much less willing to provide an incorrect answer than GPT-4-turbo–meaning GPT-4-vision-preview less confidently incorrect. In the case of when images are provided upon request from the instrument agent we find that GPT-4-turbo obtains an accuracy of 20% and GPT-4-vision-preview obtains 13% (Fig. 8). We find that images are only requested from the instrument reader in 46% of interactions, which is likely one of the leading factors behind the reduced accuracy.
@@
Previous work has made progress in the direction of clinical decision making using simulations of patients and doctors, aiming to develop AI that can diagnose through conversation. This model, titled AMIE (Articulate Medical Intelligence Explorer)[26], demonstrates improved diagnostic accuracy and performance on 28 of the 32 proposed axes from the perspective of specialist physicians and 24 of 26 axes from the perspective of patient actors. While these results are exciting for medical AI, this work remains closed-source and is not accessible for reproducibility or further studies. Additionally, this work focused only on diagnosing patients through history-taking, and did not include the ability to make decisions about which tests needed to be performed and was not configurable for multimodal clinical settings such as those with medical images or charts.
@@ Discussion
One limitation for the evaluations presented in this benchmark is that it is currently unknown what data was used to train GPT-4 and GPT-3.5. While previous works have cited GPT-4s accuracy as a valid measure[3, 13, 17], it is entirely possible that GPT-4/3.5 could have been trained on the MedQA test set giving it an unfair advantage on the task. Currently, Mixtral-8x7B[36] and Llama 2-70B-Chat[37] do not report training on the MedQA test or train set. Additionally, our results on varying the patient LLM suggest that their may be an advantage for LLMs which act as both the patient and the doctor agent, because LLMs are able to recognize their own text with high accuracy, and display disproportionate preference to that text[24].
@@ Appendix A Model details
We evaluate four language models to serve as the doctor agent (the diagnostic model): GPT-3.5, GPT-4[42], Mixtral-8x7B[36], and Llama 2 70B-chat[37]. Otherwise, for the patient, measurement, and moderator agent we use GPT-4. Briefly, we discuss the details of each model below starting with language models followed by common language models.
GPT-4 & GPT-3.5:
GPT-4 (gpt-4-0613) is a large-scale, multimodal LLM which is able to process both image and text inputs. GPT-3.5 (gpt-3.5-turbo-0613) is a subclass of GPT-3 (a 170B parameter model)[43] fine-tuned on additional tokens and with human feedback[44]. Currently, the details regarding the architecture, dataset, and training methodologies of GPT-3.5 and GPT-4 have not been not publicly disclosed. However, existing technical documentation indicates that both models are high-performing in medical and biological subjects, with GPT-4 showing superior performance compared to GPT-3.5 in knowledge assessments[42, 3].

### v2

@@ Abstract
Diagnosing and managing a patient is a complex, sequential decision making process that requires physicians to obtain information—such as which tests to perform—and to act upon it. Recent advances in artificial intelligence (AI) and large language models (LLMs) promise to profoundly impact clinical care. However, current evaluation schemes overrely on static medical question-answering benchmarks, falling short on interactive decision-making that is required in real-life clinical work. Here, we present AgentClinic: a multimodal benchmark to evaluate LLMs in their ability to operate as agents in simulated clinical environments. In our benchmark, the doctor agent must uncover the patient’s diagnosis through dialogue and active data collection. We present two open medical agent benchmarks: a multimodal image and dialogue environment, AgentClinic-NEJM, and a dialogue-only environment, AgentClinic-MedQA. We embed cognitive and implicit biases both in patient and doctor agents to emulate realistic interactions between biased agents. We find that introducing bias leads to large reductions in diagnostic accuracy of the doctor agents, as well as reduced compliance, confidence, and follow-up consultation willingness in patient agents. Evaluating a suite of state-of-the-art LLMs, we find that several models that excel in benchmarks like MedQA are performing poorly in AgentClinic-MedQA. We find that the LLM used in the patient agent is an important factor for performance in the AgentClinic benchmark. We show that both having limited interactions as well as too many interaction reduces diagnostic accuracy in doctor agents. The code and data for this work is publicly available at AgentClinic.github.io.
@@ Introduction
    Results of the diagnostic accuracy of five language models on AgentClinic-MedQA: GPT-4, GPT-4o, Mixtral-8x7B, GPT-3.5, and Llama 2 70B-chat. We also evaluate three language models on the multimodal AgentClinic-NEJM benchmark: GPT-4o, GPT-4-turbo, and GPT-4-vision-preview.
@@ Results
Here we discuss the accuracy of various language models on AgentClinic-MedQA. We evaluate five models in total: GPT-4, GPT-4o, Mixtral-8x7B, GPT-3.5, and Llama 2 70B-chat (discussed in detail in Appendix A). Each model acts as the doctor agent, attempting to diagnose the patient agent through dialogue. The doctor agent is allowed N=20 patient and measurement interactions before a diagnosis must be made. For this evaluation, we use GPT-4 as the patient agent for consistency. The accuracies of the models are presented in Figure 5: GPT-4 at 52%, GPT-4o and GPT-3.5 at 38%, Mixtral-8x7B at 37%, and Llama 2 at 70B-chat 9%.
@@
Many types of diagnoses require the physician to visually inspect the patient, such as with infections and rashes. Additionally, imaging tools such as X-ray, CT, and MRI provide a detailed and rich view into the patient, with hospitalized patients receiving an average of 1.42 diagnostic images per patient stay[25]. However, the previous experiments in this work and prior work[26] provided measurement results through text, and did not explore the ability of the model to understand visual context. Here, we evaluate three multimodal LLMs, GPT-4o (2024-05-13), GPT-4-turbo (2024-04-09) and GPT-4-vision-preview (0125), in a diagnostic settings that require interacting through both dialogue as well as understanding image readings. We collect our questions from New England Journal of Medicine (NEJM) case challenges. These published cases are presented as diagnostic challenges from real medical scenarios, and have an associated pathology-confirmed diagnosis. We curate 15 challenges from a sample of 932 total cases for AgentClinic-NEJM. While for human viewers, these cases are provided with a set of multiple choice answers, we chose to not provide these options to the doctor agent and instead keep the problems open-ended.
@@
When the image is provided initially to the doctor agent, across 15 multimodal patient settings we find that GPT-4o obtains an accuracy of 47%, whereas GPT-4-turbo and GPT-4-vision-preview obtain an accuracy of 27% (Fig. 8). We also find that for the provided incorrect responses from GPT-4-turbo, the answer that was provided was among those listed in the multiple choice options 60% of the time. Despite having the same accuracy, we find that GPT-4-vision-preview was much less willing to provide an incorrect answer than GPT-4-turbo–meaning GPT-4-vision-preview less confidently incorrect. In the case of when images are provided upon request from the instrument agent we find that GPT-4o obtains an accuracy of 27%, GPT-4-turbo obtains 20% and GPT-4-vision-preview obtains 13% (Fig. 8). We find that images are only requested from the instrument reader in 46% of interactions with GPT-4-turbo and GPT-4-vision-preview, which is likely one of the leading factors behind the reduced accuracy.
@@
Previous work has made progress in the direction of clinical decision making using simulations of patients and doctors, aiming to develop AI that can diagnose through conversation. This model, titled AMIE (Articulate Medical Intelligence Explorer)[26], demonstrates improved diagnostic accuracy and performance on 28 of the 32 proposed axes from the perspective of specialist physicians and 24 of 26 axes from the perspective of patient actors. While these results are exciting for medical AI, this work remains closed-source and is not accessible for reproducibility or further studies. Additionally, this work focused only on diagnosing patients through history-taking, and did not include the ability to make decisions about which tests needed to be performed and was not configurable for multimodal clinical settings such as those with medical images or charts. Similar to AIME, the CRAFT-MD benchmark[36] proposes evaluating LLMs through natural dialogues on dermatology questions, however without the use of images. Additionally, neither of these works demonstrate performance in the presence of bias, with multimodal input, or using a measurement agent. There has also been work which shows simulated doctor agents can improve medical QA performance through turn-based dialogue, where various medical specialist agents converse[37].
@@ Discussion
One limitation for the evaluations presented in this benchmark is that it is currently unknown what data was used to train GPT-4, GPT-4o, and GPT-3.5. While previous works have cited GPT-4s accuracy as a valid measure[3, 13, 17], it is entirely possible that GPT-4/4o/3.5 could have been trained on the MedQA test set giving it an unfair advantage on the task. Currently, Mixtral-8x7B[38] and Llama 2-70B-Chat[39] do not report training on the MedQA test or train set. Additionally, our results on varying the patient LLM suggest that their may be an advantage for LLMs which act as both the patient and the doctor agent, because LLMs are able to recognize their own text with high accuracy, and display disproportionate preference to that text[24].
@@ Appendix A Model details
We evaluate five language models to serve as the doctor agent (the diagnostic model): GPT-3.5, GPT-4[44], GPT-4o, Mixtral-8x7B[38], and Llama 2 70B-chat[39]. Otherwise, for the patient, measurement, and moderator agent we use GPT-4. Briefly, we discuss the details of each model below starting with language models followed by common language models.
GPT-4, GPT-4o, & GPT-3.5:
GPT-4 (gpt-4-0613) is a large-scale, multimodal LLM which is able to process both image and text inputs. GPT-3.5 (gpt-3.5-turbo-0613) is a subclass of GPT-3 (a 170B parameter model)[45] fine-tuned on additional tokens and with human feedback[46]. Currently, the details regarding the architecture, dataset, and training methodologies of GPT-3.5, GPT-4o (gpt-4o-2024-05-13), and GPT-4 have not been not publicly disclosed. However, existing technical documentation indicates that both models are high-performing in medical and biological subjects, with GPT-4 showing superior performance compared to GPT-3.5 in knowledge assessments[44, 3].